# Prompt-Guided Internal States for Hallucination Detection of Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities across a variety of tasks in different domains. However, they sometimes generate responses that are logically coherent but factually incorrect or misleading, which is known as LLM hallucinations. Data-driven supervised methods train hallucination detectors by leveraging the internal states of LLMs, but detectors trained on specific domains often struggle to generalize well to other domains. In this paper, we aim to enhance the cross-domain performance of supervised detectors with only in-domain data. We propose a novel framework, prompt-guided internal states for hallucination detection of LLMs, namely PRISM. By utilizing appropriate prompts to guide changes in the structure related to text truthfulness within the LLM’s internal states, we make this structure more salient and consistent across texts from different domains. We integrated our framework with existing hallucination detection methods and conducted experiments on datasets from different domains. The experimental results indicate that our framework significantly enhances the cross-domain generalization of existing hallucination detection methods [^1] We have open-sourced all the code, data, and models in GitHub: https://github.com/fujie-math/PRISM.
@@ 1 Introduction
In recent years, large language models (LLMs) have demonstrated remarkable capabilities across a variety of tasks in different domains Dinan et al. (2018); Brown (2020); Zhang et al. (2022); Chowdhery et al. (2023). However, the hallucination problem in LLMs poses a potential threat to their practical application in many scenarios. LLM hallucinations refers to the cases where LLMs generate responses that are logically coherent but factually incorrect or misleading Zhou et al. (2020); Ji et al. (2023); Su et al. (2024). These hallucinated responses may be blindly accepted, leading users to learn incorrect information or take inappropriate actions. Therefore, detecting hallucinations in LLM-generated content becomes particularly meaningful. By effectively identifying incorrect generated content through a hallucination detector, users can be alerted to verify the accuracy of certain portions of the generated content and guide the LLM to regenerate relevant sections, thus preventing potential issues.
@@
The unsupervised paradigm focuses on assessing the confidence of generated text and rejects the low-confidence ones. For example, the probability information for each token generated by LLMs can serve as a measure of hallucination Kadavath et al. (2022); Zhang et al. (2023); Quevedo et al. (2024); Hou et al. (2024). Additionally, the consistency among multiple responses generated by the LLM to the same question can also be used to evaluate their confidence Manakul et al. (2023). Chen et al. (2024) shifts the judgment of consistency among multiple responses to their corresponding internal vectors. However, these methods often struggle to achieve ideal detection accuracy or require a significant amount of additional response time Su et al. (2024).
@@
For this reason, researchers have begun exploring the use of data-driven supervised methods for detection. These methods are generally based on the assumption that LLMs can recognize they have generated incorrect content, which is reflected in specific patterns within their internal states. Marks and Tegmark (2023) reveal that true and false statements have distinguishable geometric structures within the LLM, allowing us to build classifiers by learning this structure. Marks and Tegmark (2023) and Sky et al. (2024) utilize linear structures to distinguish between different statements. Azaria and Mitchell (2023) train fully connected networks to act as hallucination detectors by using activation values within the LLM.
@@
Supervised detectors trained on specific domains often struggle to achieve good generalization performance in other domains. For example, Bürger et al. (2024) found that the structural differences in the LLM’s internal states for affirmative sentences differ significantly from those for negated sentences when distinguishing between true and false statements. To address this issue, many studies focus on constructing diverse datasets or performing feature selection within the LLM to achieve better generalization performance Chen et al. (2023); Bürger et al. (2024); Liu et al. (2024). However, these methods require collecting additional training data from other domains, which is resource-intensive. In this paper, we aim to answer the following research question:
@@
Can we enhance the cross-domain performance of supervised detectors without using datasets from other domains?
@@
Driven by this research question, we propose a novel framework called PRISM, which stands for prompt-guided internal states for hallucination detection of LLMs. By utilizing appropriate prompts to guide changes in the structure related to text truthfulness within the LLM’s internal states, we make this structure more salient and consistent across texts from different domains. This enables detectors trained in one domain to perform well in others without additional data. Our approach is based on the insight that while the internal states of LLMs encode rich semantic information, they are primarily optimized during pre-training to predict the next token rather than for hallucination detection. As a result, the directly extracted internal states contain a lot of domain-specific information unrelated to truthfulness, leading to detectors that are specific to certain domains and unable to achieve good cross-domain generalization.
@@
Due to LLM’s powerful ability to understand and follow instructions, we explore the use of prompt-guided methods to generate internal states more focused on truthfulness. We employed various methods to investigate the effect of prompts from different perspectives. The findings indicate that the introduction of appropriate prompts can significantly improve the salience of the structure related to text truthfulness in the LLM’s internal states and make this structure more consistent across different domain datasets. We also provide a simple and effective method to generate and select appropriate prompts for hallucination detection tasks. By combining the prompt template with the text to be evaluated and inputting them into the LLM, we obtain internal states that are better suited as features for hallucination detection tasks. Subsequently, we can integrate with existing hallucination detection methods to construct suitable detectors for classifying the text.
@@
In summary, the contributions of this paper are as follows: (1) We conduct an in-depth investigation into how specific prompts influence the internal state structures of LLMs with regard to text truthfulness. (2) We propose PRISM, a novel framework that utilizes prompt-guided internal states to enhance hallucination detection in LLMs. (3) Successfully integrated PRISM with existing hallucination detection methodologies, demonstrating its ability to significantly enhance generalization performance across various domains.
@@
Figure 1: Visualization of the LLaMA2-7B-Chat model’s internal states by using 2-dimensional PCA on the True-False dataset before and after the introduction of Prompt 1, where label 0 represents false statements and label 1 represents true statements. Additionally, a logistic regression model is fitted to distinguish between true and false statements, with the decision boundary shown as a black dashed line.
2 Datasets
@@
Another dataset we used is from Bürger et al. (2024) and consists of 24 sub-datasets with 6 different topics: "animal_class," "cities," "inventors," "element_symb," "facts," and "sp_en_trans," as well as 4 different grammatical structures: affirmative statements, negated statements, logical conjunctions, and logical disjunctions. Among the affirmative statements, "element_symb," "animal_class," "inventors," and "facts" are subsets of the True-False datasets, while the other sub-datasets were constructed based on the datasets used by Marks and Tegmark (2023) and Levinstein and Herrmann (2024).
Due to data imbalance across topics, we focus more on grammatical structure variations within this dataset. Affirmative statements are structured similarly to the examples in the True-False dataset, while negated statements are formed by negating the affirmative statements using the word "not." The sentences in logical conjunctions and logical disjunctions were constructed by sampling and combining sentences from the affirmative statements and then connecting them with "and" or "or." The sampling rules were carefully designed to ensure the overall dataset is balanced between true and false statements. For clarity, we refer to this dataset as LogicStruct throughout this paper and present additional information about it in Appendix A.
@@ 3 Prompt-Guided Internal States
Several previous studies have focused on leveraging the internal states of LLMs for hallucination detection Azaria and Mitchell (2023); Liu et al. (2024); Su et al. (2024). They assume that LLMs can recognize they have generated incorrect content, which is reflected in specific patterns within their internal states. Therefore, these studies typically select contextualized embeddings corresponding to specific tokens in certain hidden layers of LLMs, and then use them as features to train hallucination detectors. However, these embeddings were originally intended to guide LLMs’ text generation and they encode various information of the related text. As a result, the information related to text truthfulness becomes intertwined with domain-specific details, making it difficult for the detector to identify a consistent structure related to text truthfulness in LLMs’ internal states.
Therefore, we hope to find a way to guide changes in the LLM’s internal states such that the structure related to text truthfulness becomes more salient and consistent across texts from different domains. This would help the detector learn this structure and enhance its generalization. To achieve this goal, we experimented with a simple prompt:
@@
This prompt directly asks the LLM the truthfulness of a specific statement. When we input it into the LLM, the generation of new tokens will revolve around this question, which will be reflected in the LLM’s internal states. Although the prompt itself cannot be classified as true or false, it contains statements with different truthfulness, so this difference will be still reflected in the internal states of the LLM, resulting in a new structure related to text truthfulness.
In the following two subsections, we will use various methods to demonstrate the changes in the structure related to text truthfulness within the LLM’s internal states before and after the introduction of the prompt and explain how these changes will effect the hallucination detection task. Our analysis is conducted on the sub-datasets of the True-False dataset. The feature vectors used in our analysis are the embeddings corresponding to the last token in the last layer of LLaMA2-7B-Chat. More experimental results under other conditions are presented in Section 5.
@@
Figure 2: Comparison of variance ratios before and after the introduction of Prompt 1 on the True-False dataset.
Table 1: Cosine similarity between truthfulness directions across different topics in the True-False dataset, before and after Prompt 1 addition.
@@
3.1 Effect of the Prompt on Structure Salience
In this subsection, we apply principal component analysis and variance analysis in specific directions to reveal that the introduction of the prompt can make the structure related to text truthfulness within the LLM’s internal states more salient.
Some previous studies Marks and Tegmark (2023); Bürger et al. (2024) have shown that true and false statements have distinguishable geometric structures within the LLM on certain datasets, and these structures can be directly observed after applying Principal Component Analysis to reduce the dimensionality of the relevant data. Therefore, we first apply PCA to observe the salience of the structure related to text truthfulness before and after the introduction of Prompt 1.
From Figure 1, we can clearly observe that in all six datasets, the introduction of the prompt enables the first two principal components to effectively separate the true and false statements. Since the first two principal components represent the directions with the largest variance among the embeddings, this indicates that the introduction of the prompt makes the structure related to text truthfulness in the embeddings more salient. However, this is just an intuitive but imprecise perspective, as we cannot be certain about the actual meaning represented by the first two principal components.
To accurately describe the differences in structure salience, we conduct a more detailed variance analysis. We believe that the proportion of variance along a particular direction relative to the total variance can reflect the salient of the structure in that direction. The larger this proportion, the more salient the relevant structure is compared to others. Therefore, to measure the salience of the structure related to text truthfulness, we need to identify a direction in the vector space that is associated with this structure.
A simple method is to subtract the average embedding of false statements from the average embedding of true statements within the dataset. We can refer to this direction as the truthfulness direction. This direction was used by Marks and Tegmark (2023) for classifying statements with different truthfulness. Its specific formula is as follows:
@@ 3.1 Effect of the Prompt on Structure Salience
where $\theta$ represents the truthfulness direction in dataset $D$, $v_{i}^{+}$ and $v_{i}^{-}$ denote the embeddings corresponding to true and false statements in dataset $D$, respectively, and $N^{+}$ and $N^{-}$ represent their counts.
Then we can perform variance analysis along this direction. Let $v_{i}$ represents the embedding corresponding to a statement within dataset $D$, and let $N$ denote the total number of statements. By calculating
@@
we can use the de-centered vectors $\hat{v_{i}}$ to form the data matrix $X^{T}=(\hat{v_{1}},\hat{v_{2}},\dots,\hat{v_{N}})$. Next, we compute the covariance matrix corresponding to the data matrix:
@@
and the variance along the truth direction:
@@
to represent the salience of the structure related to text truthfulness. We calculate the corresponding variance ratios on the six sub-datasets of the True-False dataset and presented the results in Figure 2.
We can see that after the introduction of the prompt, this ratio shows a significant increase for all sub-datasets, indicating that the differences among the data begin to concentrate more along the truthfulness direction. This enhances the salience of the structure related to text truthfulness in these embeddings, making it easier for the hallucination detector to learn.
3.2 Effect of the Prompt on Structural Consistency
In this subsection, we focus on analyzing the consistency of structures related to text truthfulness across different datasets. We calculated the cosine similarity between truthfulness directions and observed that the introduction of the prompt makes these structures more consistent.
The salience of the structure related to text truthfulness may impact the training performance of the hallucination detector. However, when we aim to apply a detector trained in one domain to others, the consistency of this structure becomes particularly important. A more consistent structure will lead to better generalization performance. Therefore, we continue to use the ’truthfulness direction’ defined by Equation (1) to analyze the consistency of this structure across different datasets. We use cosine similarity to represent this consistency:
@@ 3.2 Effect of the Prompt on Structural Consistency
where $c_{ij}$ represents the cosine of the angle between the truthfulness directions of datasets $D_{i}$ and $D_{j}$.
The calculation results for the six sub-datasets of the True-False dataset are presented in Table 1. We can see that the introduction of the prompt significantly increases the cosine similarity between the truthfulness directions across datasets, indicating that the truthfulness directions between different datasets become more consistent. This consistency will help improve the generalization of the related detector, which uses the LLM’s internal states as input features to determine the truthfulness of statements.
Table 2: Average variance ratios for 10 candidate prompt templates from Appendix C along with their ranking.
Prompts I II III IV V VI VII VIII IX X Without

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities across a variety of tasks in different domains. However, they sometimes generate responses that are logically coherent but factually incorrect or misleading, which is known as LLM hallucinations. Data-driven supervised methods train hallucination detectors by leveraging the internal states of LLMs, but detectors trained on specific domains often struggle to generalize well to other domains. In this paper, we aim to enhance the cross-domain performance of supervised detectors with only in-domain data. We propose a novel framework, prompt-guided internal states for hallucination detection of LLMs, namely PRISM. By utilizing appropriate prompts to guide changes to the structure related to text truthfulness in LLMs’ internal states, we make this structure more salient and consistent across texts from different domains. We integrated our framework with existing hallucination detection methods and conducted experiments on datasets from different domains. The experimental results indicate that our framework significantly enhances the cross-domain generalization of existing hallucination detection methods [^1] We have open-sourced all the code and data in GitHub: https://github.com/fujie-math/PRISM.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have demonstrated remarkable capabilities across a variety of tasks in different domains Dinan et al. (2019); Brown et al. (2020); Zhang et al. (2022); Chowdhery et al. (2023); Touvron et al. (2023a). However, the hallucination problem in LLMs poses a potential threat to their practical application in many scenarios. LLM hallucinations refer to the cases where LLMs generate responses that are logically coherent but factually incorrect or misleading Zhou et al. (2021); Ji et al. (2023); Su et al. (2024). These hallucinated responses may be blindly accepted, leading users to learn incorrect information or take inappropriate actions. Therefore, detecting hallucinations in LLM-generated content becomes particularly meaningful. By using hallucination detectors to help identify incorrect generated content, users can be alerted to verify the accuracy of LLMs’ responses, thus preventing potential issues.
@@
The unsupervised paradigm focuses on assessing the confidence of LLM-generated content and rejects the low-confidence outputs. For example, the probability information of each token generated by LLMs can serve as a measure of hallucination Kadavath et al. (2022); Zhang et al. (2023); Quevedo et al. (2024); Hou et al. (2024). Additionally, the consistency among multiple responses generated by LLMs to the same question can also be used to evaluate their confidence Manakul et al. (2023). Furthermore, Chen et al. (2024) shifts the consistency judgment from multiple responses to their corresponding activation values in LLMs. However, these methods often struggle to achieve ideal detection accuracy or require a significant amount of additional response time Su et al. (2024).
@@
For this reason, researchers have begun exploring the use of data-driven supervised methods for hallucination detection. These methods are generally based on the assumption that LLMs can recognize they have generated incorrect content, which is reflected in specific patterns in their internal states. Marks and Tegmark (2023) reveal that true and false statements have discernible geometric structures in LLMs’ internal states, allowing us to build classifiers by learning this structure. Marks and Tegmark (2023) and CH-Wang et al. (2024) utilize linear structures to distinguish between different statements, while Azaria and Mitchell (2023) train neural networks to act as hallucination detectors.
@@
Supervised detectors trained on specific domains often struggle to achieve good generalization performance in other domains. For example, Bürger et al. (2024) found that when distinguishing between true and false statements, the structural differences in LLMs’ internal states are significantly different for affirmative and negated sentences. To address this issue, many studies focus on constructing diverse datasets or performing feature selection within LLMs’ internal states to achieve better generalization performance Chen et al. (2023); Bürger et al. (2024); Liu et al. (2024). However, these methods require collecting additional training data from other domains, which is resource-intensive. In this paper, we aim to answer the following question:
@@
Can we enhance the cross-domain performance of supervised detectors with only in-domain data?
@@
Driven by this research question, we propose a novel framework called PRISM, which stands for prompt-guided internal states for hallucination detection of LLMs. By utilizing appropriate prompts to guide changes to the structure related to text truthfulness in LLMs’ internal states, we make this structure more salient and consistent across texts from different domains, which enables detectors trained in one domain to also perform well in others without additional data. Our approach is based on the insight that while LLMs’ internal states encode rich semantic information, they are primarily optimized during pre-training to predict the next token rather than for hallucination detection. As a result, the directly extracted internal states contain lots of domain-specific information that is not related to text truthfulness, leading to detectors that are specific to certain domains and unable to achieve good cross-domain generalization performance.
@@
Due to the powerful ability of LLMs to understand and follow instructions, we explore the use of prompt-guided methods to generate internal states more focused on text truthfulness. We employed various methods to investigate the effect of prompts from different perspectives. The findings indicate that the introduction of appropriate prompts can significantly improve the salience of the structure related to text truthfulness in LLMs’ internal states and make this structure more consistent across different domain datasets. We also provide a simple and effective method to generate and select appropriate prompts for hallucination detection tasks. By combining prompt templates with the text to be evaluated and inputting them into LLMs, we can obtain internal states that are better suited as features for hallucination detection tasks. Subsequently, we can integrate the prompt-guided internal states with existing hallucination detection methods to construct more advanced detectors.
@@
In summary, the contributions of this paper are as follows: (1) We conduct an in-depth investigation into how specific prompts guide the change to the structure related to text truthfulness in LLMs’ internal states. (2) We propose PRISM, a novel framework that utilizes prompt-guided internal states to enhance hallucination detection of LLMs. (3) We integrate PRISM with existing hallucination detection methodologies, demonstrating its ability to significantly enhance the generalization performance of detectors across different domains.
@@
Figure 1: Visualization of the LLaMA2-7B-Chat model’s internal states by using the 2-dimensional PCA on the True-False dataset before and after the introduction of Prompt 1, where label 0 represents false statements and label 1 represents true statements. Additionally, a logistic regression model was fitted to distinguish between true and false statements, with the decision boundary shown as a black dashed line.
2 Preliminary
@@
Another dataset we used is from Bürger et al. (2024) and consists of 24 sub-datasets with 6 different topics: "animal_class," "cities," "inventors," "element_symb," "facts," and "sp_en_trans," as well as 4 different grammatical structures: affirmative statements, negated statements, logical conjunctions, and logical disjunctions. Affirmative statements were structured similarly to the examples in the True-False dataset, while negated statements were formed by negating the affirmative statements using the word "not." The sentences in logical conjunctions and logical disjunctions were constructed by sampling sentences from the affirmative statements and then connecting them with "and" or "or." The number of true and false statements within each grammatical structure is balanced. For clarity, we refer to this dataset as LogicStruct throughout this paper and present additional information about it in Appendix A.
@@ 3 Prompt-Guided Internal States
Several previous studies have focused on leveraging LLMs’ internal states for hallucination detection Azaria and Mitchell (2023); Liu et al. (2024); Su et al. (2024). They assume that LLMs can recognize they have generated incorrect content, which is reflected in specific patterns in their internal states. Therefore, these studies typically select contextualized embeddings corresponding to specific tokens in certain layers of LLMs, and then use them as features to train hallucination detectors. However, these embeddings were originally intended to guide text generation and they encode various information of the related text. As a result, the information related to text truthfulness becomes intertwined with domain-specific details, making it difficult for detectors to identify a consistent structure related to text truthfulness in LLMs’ internal states.
Therefore, we hope to guide changes in LLMs’ internal states such that the structure related to text truthfulness becomes more salient and consistent across texts from different domains. This would help detectors learn this structure and enhance their generalization performance. To achieve this goal, we experimented with a simple prompt:
@@
This prompt directly asks LLMs the truthfulness of specific statements. When we input it into LLMs, the generation of new tokens will revolve around this question, which will be reflected in LLMs’ internal states.
In the following two subsections, we will use various methods to demonstrate the changes in LLMs’ internal states before and after the introduction of Prompt 1 and explain how these changes will effect the hallucination detection task. Our analysis is conducted on the six sub-datasets of the True-False dataset. The feature vectors we used are the embeddings corresponding to the last token in the final layer of LLaMA2-7B-Chat.
3.1 Effect of Prompt on Structural Salience
Some previous studies Marks and Tegmark (2023); Bürger et al. (2024) have shown that, on certain datasets, true and false statements have distinguishable geometric structures in LLMs’ internal states and these structures can be directly observed after applying Principal Component Analysis (PCA) to reduce the dimensionality of relevant data. Therefore, we utilized the PCA to observe the salience of the structure related to text truthfulness before and after the introduction of Prompt 1.
From Figure 1, we can clearly observe that in all six datasets, the introduction of the prompt enables the first two principal components to effectively distinguish true and false statements. Since the first two principal components represent the directions with the largest variance among the embeddings, this indicates that the introduction of the prompt makes the structure related to text truthfulness in the embeddings more salient.
@@ 3.1 Effect of Prompt on Structural Salience
Figure 2: Comparison of the variance ratios before and after the introduction of Prompt 1 on each sub-dataset of the True-False dataset.
Moreover, to accurately describe the differences in structural salience, we performed a more detailed variance analysis. We believe that the proportion of the variance along a particular direction relative to the total variance can reflect the salience of the structure in that direction. Therefore, we subtracted the average embedding of false statements from the average embedding of true statements within each dataset, and referred to this direction as the ’truthfulness direction’. This was first used by Marks and Tegmark (2023) and its formula is as follows:
Table 1: The cosine similarity between the truthfulness directions across different topics in the True-False dataset, before and after Prompt 1 addition.
@@
where $\theta$ represents the truthfulness direction in the dataset $D$, $v_{i}^{+}$ and $v_{i}^{-}$ denote the embeddings corresponding to true and false statements, respectively, and $N^{+}$ and $N^{-}$ represent their counts.
Then we performed variance analysis along this direction. Let $v_{i}$ represent the embedding corresponding to a specific statement within dataset $D$, and let $N$ denote the total number of statements. By calculating
@@
we can use the de-centered vectors $\hat{v_{i}}$ to form a data matrix $X=(\hat{v_{1}},\hat{v_{2}},\dots,\hat{v_{N}})^{T}$, and then compute the covariance matrix corresponding to $X$:
@@
and the variance along the truthfulness direction:
@@
to represent the salience of the structure related to text truthfulness in dataset $D$. We calculated the corresponding variance ratios on the six sub-datasets of the True-False dataset and presented the results in Figure 2.
We can observe that after the introduction of the prompt, this ratio shows a significant increase for all sub-datasets, indicating that the differences among the embeddings begin to concentrate more along the truthfulness direction. This enhances the salience of the structure related to text truthfulness in LLMs’ internal states, making it easier for hallucination detectors to learn.
3.2 Effect of Prompt on Structural Consistency
When we aim to apply detectors trained in one domain to others, the consistency of this structure becomes particularly important. A more consistent structure will lead to better generalization performance. Therefore, we continued to utilize the ’truthfulness direction’ defined by Equation (1) to analyze the consistency of this structure across different datasets and use the cosine similarity to represent this consistency:
@@ 3.2 Effect of Prompt on Structural Consistency
where $c_{ij}$ represents the cosine of the angle between two truthfulness directions.
The calculation results for the six sub-datasets of the True-False dataset are presented in Table 1. We can see that the introduction of the prompt significantly increases the cosine similarity between the truthfulness directions across different datasets, indicating that the structure related to text truthfulness between different datasets becomes more consistent. This consistency will help improve the generalization performance of related detectors, which use LLMs’ internal states as input features to determine the truthfulness of statements.
4 Methodology
The analyses in sections 3 indicate that the introduction of appropriate prompts can guide changes to the structure related to text truthfulness in LLMs’ internal states, and these changes will facilitate our use of LLMs’ internal states as features to perform the hallucination detection task. Based on this insight, we propose a novel framework PRISM, i.e., PRompt-guided Internal States for hallucination detection of Large Language Models.
Our framework consists of two components:
First, we generate a large number of candidate prompt templates and select the one that best fits the current task. We manually construct a prompt template $P$ related to the hallucination detection task, and then we use the language model $L$ to generate multiple candidate prompt templates $\{P_{i}\}_{i=1,\dots,N}$ that are similar in meaning but different in structure. For each $P_{i}$, we combine it with the text $a_{j}$ from the labeled dataset $A$ to obtain the text $P_{i}(a_{j})$, and then input it into $L$ to obtain the corresponding feature vector $v_{ij}$ in the internal states. Subsequently, we use all feature vectors $\{v_{ij}\}_{j=1,\dots,M}$ to calculate the structure salience index $R_{i}$ corresponding to $P_{i}$ on $A$ and select the prompt template with the highest index as our final prompt template $P_{s}$.
Next, we integrate the prompt-guided internal states with existing hallucination detection methods to construct a more advanced detector. We train the hallucination detector $D$ on the dataset $A$ using all feature vectors $\{v_{sj}\}_{j=1,\dots,M}$ and their labels. For new text $T$ to be detected, we similarly combine it with the prompt template $P_{s}$ to obtain the text $P_{s}(T)$ and input it into the language model $L$ to obtain the corresponding feature vector $v$. Finally, we input $v$ into the hallucination detector $D$ to obtain its hallucination label $H$.
In the following subsections, we will provide a more detailed explanation of some key steps in our framework.
Table 2: Average variance ratios for the 10 candidate prompt templates along with their ranking.
Prompt I II III IV V VI VII VIII IX X Without